# Clone Deterministic 3D Worlds

### v2

@@ Abstract
A world model is an internal model that simulates how the world evolves. Given past observations and actions, it predicts the future physical state of both the embodied agent and its environment. Accurate world models are essential for enabling agents to think, plan, and reason effectively in complex, dynamic settings. However, existing world models often focus on random generation of open worlds, but neglect the need for high-fidelity modeling of deterministic scenarios (such as fixed-map mazes and static space robot navigation). In this work, we take a step toward building a truly accurate world model by addressing a fundamental yet open problem: constructing a model that can fully clone a deterministic 3D world. 1) Through diagnostic experiment, we quantitatively demonstrate that high-fidelity cloning is feasible and the primary bottleneck for long-horizon fidelity is the geometric structure of the latent representation, not the dynamics model itself. 2) Building on this insight, we show that applying temporal contrastive learning principle as a geometric regularization can effectively curate a latent space that better reflects the underlying physical state manifold, demonstrating that contrastive constraints can serve as a powerful inductive bias for stable world modeling; we call this approach Geometrically-Regularized World Models (GRWM). At its core is a lightweight geometric regularization module that can be seamlessly integrated into standard autoencoders, reshaping their latent space to provide a stable foundation for effective dynamics modeling. By focusing on representation quality, GRWM offers a simple yet powerful pipeline for improving world model fidelity.
@@ 1 Introduction
Creating a high-fidelity, interactive clone of an environment purely from observational data has long been a central ambition in artificial intelligence. Such a clone can serve as a simulator for reinforcement learning agents (hafner2020dreamer; hao2025neural; sutton), enable task planning in robotics (mendonca2023structured; unitree), and facilitate controllable content generation in games (alonso2024diffusionworldmodelingvisual; oasis2024; bruce2024genie). World models are the primary tools for achieving this goal: they aim to capture an environment’s dynamics, predict its future states, and simulate its evolution (dyna; ha2018world; schrittwieser2020mastering; lecun2022path).
Most current world models focus on open-world settings (oasis2024; he2025matrix; genie3; sora), generating unconstrained and often random environments in which each simulation produces a different world. This approach is intended to enhance generalization by exposing agents to diverse training scenarios (jang2025dreamgen; alonso2024diffusionworldmodelingvisual; hafner2020dreamer; hafner2019learning; bruce2024genie). However, the emphasis on randomness can lead to unstable dynamics, making such models ill-suited for applications that demand reliable prediction and precise planning in fixed tasks. In these cases, they often produce futures that are merely plausible, not faithful.
@@
  • Representation learning. Exteroceptive sensory data, such as images, are high-dimensional and encode complex, nonlinear mappings from underlying physical processes (battaglia2018relational; wang2020understanding; tschannen2018recent; arats2021improving; kang2024far). This makes accurate future-state prediction difficult even under full observability (ha2018world; hafner2019learning). For example, a simple spatial translation in the physical world can correspond to a highly nonlinear trajectory in pixel space (tancik2020fourier; nonlinear1; nonlinear2). Addressing this requires a representation space that faithfully encodes the underlying physical states while minimizing information loss and noise, thereby simplifying the subsequent dynamics modeling task (wang2022denoised; fang2024towards; seo2023masked; karlsson2023predictive; battaglia2018relational; zhou2024dinowmworldmodelspretrained).
@@
  • Dynamics modeling. Even with an optimal representation, the dynamics model must capture a broad range of transition patterns — including 3D transformations, logical rules, causal dependencies, and temporal memory (dynamics1; dynamics2; dynamics3; dynamics4; hao2025neural; chen2024diffusion; huang2025selfforcing). The challenge lies in building a unified mechanism that can accurately model all these regularities (dynamics1; dynamics4; dynamics5; dynamics6).
@@
These challenges are inherently coupled: a poor representation forces the dynamics model to operate in a noisy, entangled latent space, increasing prediction complexity and reducing generalization; conversely, a well-structured representation is of limited utility if the dynamics model cannot capture the full spectrum of transition patterns (representation1; representation2; tschannen2018recent; blau2019rethinking; kouzelis2025eq; stab). Progress toward generic neural world simulation therefore demands a co-design of representation learning and dynamics modeling, ensuring that the latent space is both physically meaningful and optimally aligned with the predictive capabilities of the dynamics model (ha2018world; seo2023masked; hafner2020dreamer; dynamics5; chen2024deep).
@@
In this work, we focus on the faithful cloning of deterministic 3D environments — settings governed by fixed rules, such as a 3D maze with a static map. We target these environments for three reasons: (1) Their consistent dynamics make them amenable to precise predictive modeling. (2) The absence of stochasticity eliminates uncertainty, enabling rigorous evaluation of a world model’s fidelity. (3) Many important real-world applications — including robot navigation in fixed spaces and game AI operating on static maps — inherently involve deterministic environments. In such cases, the objective is not to produce a plausible world, but to reproduce the unique, true trajectory of the environment. Ultimately, our goal is to construct a digital twin indistinguishable from the original in both rules and behavior.
@@
Yet, we find that achieving accurate long-horizon cloning of even simple deterministic 3D environments remains an open challenge. Across all state-of-the-art baselines we evaluated, none were able to maintain fidelity over extended horizons: small prediction errors accumulate rapidly, causing trajectories to diverge from reality after only a few steps. In contrast, when the dynamics model is provided with the environment’s underlying physical states [^1] We refer to this version as the oracle model, since it has access to the true underlying states of the environment. — rather than high-dimensional exteroceptive signals such as images — it can produce remarkably accurate long-horizon predictions. This observation suggests that the effectiveness of a world model is fundamentally constrained by the structure of its latent representation space, and thus raises a natural question: How can we build a self-supervised representation that aligns with the underlying physical states, enabling stable and accurate long-horizon predictions?
@@
A natural next step is to investigate how to bridge the gap to oracle-level performance through unsupervised representation learning. We show that applying the temporal contrastive learning principle (foldiak1991learning; wiskott2002slow; cltt3; cltt5; cltt7; klindt2020towards) as a geometric regularization can effectively curate a latent space that better reflects the underlying physical state manifold, demonstrating that contrastive constraints can serve as a powerful inductive bias for stable world modeling; we call this approach Geometrically-Regularized World Models (GRWM). At its core is a lightweight geometric regularization module that can be seamlessly integrated into standard autoencoders, reshaping their latent space to better align with the underlying physical states and provide a stable foundation for effective dynamics modeling. Through extensive experiments, we demonstrate that improved representation geometry directly translates into higher long-horizon fidelity across multiple world models, validating the “representation matters” hypothesis. GRWM acts as a plug-and-play component that systematically unlocks long-horizon fidelity for state-of-the-art world models.
@@ 2 Related Work
Video world models aim to predict future visual observations of an environment conditioned on agent actions, often using latent generative modeling as a central approach (oasis2024; he2025matrix; parkerholder2024genie2). Latent generative world models (seo2023masked; high; ha2018world) first encode observations into a latent space via an autoencoder and then model transitions in this space to predict future frames. This paradigm is widely applicable (oasis2024; parkerholder2024genie2; genie3; bruce2024genie; sora) and has inspired recent works exploring various enhancements, including interactive controllability (i1; oasis2024; a2), architectural innovations for causality (yin2025causvid; chen2024diffusion) and persistence (framepack; a1; far; framepackv1; song2025history), and improved physical accuracy (p1; p2; p3). Despite these advances, the predominant focus has been on improving dynamics modeling, e.g., designing more expressive latent transition models (t1; huang2025selfforcing; song2025history; a1) and adding causal and relational inductive biases (mendonca2023structured; hao2025neural; chen2024diffusion). The autoencoder’s role is largely engineering: to provide a compressive latent embedding for the dynamics model [^2] Commonly, the autoencoder is trained with a small KL weight and additional perceptual or adversarial losses (high)., with most work focusing on improving reconstruction quality rather than shaping the latent geometry for stable long-horizon prediction.
Moreover, most video world models target open-world or stochastic environments, emphasizing diversity and plausibility over faithful reproduction of a unique trajectory (oasis2024; he2025matrix; genie3; bruce2024genie; sora). In contrast, deterministic environments, where the goal is to precisely clone the true sequence of states, have received less attention. Our work highlights that in such settings, representation quality can be the primary bottleneck: even strong dynamics models fail if the latent space is poorly structured, motivating methods like GRWM that explicitly regularize latent geometry.
@@
Pure reconstruction objectives often yield degenerate representations that ignore temporal cues. Contrastive learning principles (drlim; oord2018representation; chen2018sparse; chen2020simple; he2020momentum; wang2020understanding; yeh2022decoupled; garrido2022duality; wang2024pose) suggest a remedy: enforce similarity for related samples and repulsion for unrelated ones. Several methods extend this to temporal data, such as Contrastive Learning Through Time (CLTT) (cltt1; cltt2; cltt3; cltt4; cltt5; cltt6; cltt7; cltt8; wang2024pose), where contrastive objectives are used to capture temporal coherence in sequential observations.
@@ 3 Preliminary
The Next-State Prediction Problem. World models aim to perform next-state prediction: learning a function that predicts future observations given the current state and an action (BrysonHo69; ha2018world; sutton). An environment evolves through latent states $s_{t}$, which yield observations $o_{t}$ after actions $a_{t}$.In the partially observable case, $o_{t}$ gives only partial information about $s_{t}$, requiring integration of past history to form a belief state. In the fully observable case, $s_{t}\approx o_{t}$, yet predicting in observation space remains difficult because simple physical transitions (e.g., translation) correspond to complex, non-linear changes in high-dimensional pixel space. This necessitates a learned representation space that can linearize these dynamics.
@@
We adopt a latent generative framework (seo2023masked; high; ha2018world) consisting of two stages: (1) an autoencoder that learns a compressed latent representation (kingma2013auto; higgins2017beta; vqvae), and (2) a generative model that captures the dynamics in that latent space (ha2018world; hafner2019learning; bruce2024genie). Our work focuses on the representation stage, as its quality critically determines the downstream rollout performance.
@@ 4 Representation as the Bottleneck
Figure 1: Representation quality is the primary bottleneck for world model fidelity. Frame-wise MSE on the Maze 3x3 dataset. (Left) An oracle model using ground-truth states (black dotted) achieves near-zero error, establishing a performance upper bound. In contrast, a standard VAE-based world model (blue dashed) accumulates error rapidly. Our GRWM (green solid) significantly closes this gap by learning a more structurally aligned latent space (Right Bottom), while the VAE’s representation remains disorganized (Right Top). For further details, see Section 6.
We first conduct a foundational diagnostic experiment on the Maze $\times3$ environment to identify the source of error in high-fidelity cloning. We construct an “oracle” world model by training directly on the ground-truth underlying states, i.e., the agent’s coordinates $(x,y,\theta)$. This oracle, trained with full supervision, is designed to completely bypass the challenge of perception. It consists of three components: (1) An encoder that maps a history of observations to the previous state: $\hat{s}_{t-1}=f_{\text{enc}}(o_{t-k:t-1}),$ where $s_{t}=(x_{t},y_{t},\theta_{t})$. (2) A dynamics model that predicts the current state from the previous state and action: $\hat{s}_{t}=f_{\text{dyn}}(\hat{s}_{t-1},a_{t}).$ (3) A decoder that generates the current observation $\hat{o}_{t}$ from the current latent state: $\hat{o}_{t}=f_{\text{dec}}(\hat{s}_{t}).$
As shown in Figure 1, this oracle predicts future states with near-perfect accuracy. This result is a key finding: it quantitatively proves two critical points: (1) high-fidelity cloning is entirely feasible, and (2) the dynamics model itself is not the bottleneck. The challenge is perception. This oracle’s success establishes a clear performance upper bound and highlights that discovering these underlying states is a critical path toward perfect world modeling.
However, this oracle’s success contrasts with standard methods. When the exact same dynamics model is trained on the latent space of a standard VAE, its performance collapses. This massive performance gap highlights that the primary bottleneck is not the dynamics model, but the representation itself. The VAE’s latent space, optimized solely for reconstruction, is not structurally aligned with the environment’s state manifold, leading to catastrophic errors.
As a preview, Figure 1 also shows that a regularized VAE, trained with the geometric priors we introduce next, significantly closes this performance gap. This experiment solidifies our central thesis: representation quality is the primary bottleneck for world model fidelity.
@@ 5 Geometrically-Regularized World Models
We next explore how to reduce the gap to oracle-level performance through unsupervised representation learning. Our approach applies the temporal contrastive learning principle (cltt3; cltt5; cltt7; klindt2020towards) as a geometric regularizer on the autoencoder of a world model. This regularizer shapes the latent space so that its geometry better matches the underlying state manifold, providing a strong inductive bias for stable long-horizon prediction.
@@ 5.1 Temporal Contextualize Architecture
Learning underlying states from a single frame is difficult due to perceptual aliasing (alias1; alias2; alias3; alias4), where distinct states can yield nearly identical observations. For example, two different positions in a maze may appear visually indistinguishable. This also clarifies a key difference in the oracle experiment: the oracle uses coordinates, which encode the global state directly, whereas images provide only local information. A single image cannot specify the global position in environments with heavy aliasing. To recover the true state, the model must aggregate information over time, since a sequence of observations can form a unique signature of the underlying global state (cltt3; cltt5; cltt7).
@@
Each frame is first encoded independently with a 2D CNN. The resulting frame-level features are then aggregated by a causal Transformer (vaswani2017attention) with a sliding temporal window, ensuring that $z_{t}$ only attends to a limited range of past frames up to and including time $t$. This windowed design captures short-term temporal context while maintaining causality.

### v3

# Cloning Deterministic Worlds: The Critical Role of Latent Geometry in Long-Horizon World Models

@@ Abstract
A world model is an internal model that simulates how the world evolves. Given past observations and actions, it predicts the future physical state of both the embodied agent and its environment. Accurate world models are essential for enabling agents to think, plan, and reason effectively in complex, dynamic settings. However, existing world models often focus on random generation of open worlds, but neglect the need for high-fidelity modeling of deterministic scenarios (such as fixed-map mazes and static space robot navigation). In this work, we take a step toward building a truly accurate world model by addressing a fundamental yet open problem: constructing a model that can fully clone a deterministic world. 1) Through diagnostic experiment, we quantitatively demonstrate that high-fidelity cloning is feasible and the primary bottleneck for long-horizon fidelity is the geometric structure of the latent representation, not the dynamics model itself. 2) Building on this insight, we show that applying temporal contrastive learning principle as a geometric regularization can effectively curate a latent space that better reflects the underlying physical state manifold, demonstrating that contrastive constraints can serve as a powerful inductive bias for stable world modeling; we call this approach Geometrically-Regularized World Models (GRWM). At its core is a lightweight geometric regularization module that can be seamlessly integrated into standard autoencoders, reshaping their latent space to provide a stable foundation for effective dynamics modeling. By focusing on representation quality, GRWM offers a simple yet powerful pipeline for improving world model fidelity.
footnotetext: Corresponding author. Project page: https://xiafire.github.io/grwm-project-page
@@ 1 Introduction
Creating a high-fidelity, interactive clone of an environment purely from observational data has long been a central ambition in artificial intelligence. Such a clone can serve as a simulator for reinforcement learning agents [32, 33, 67], enable task planning in robotics [55, 71], and facilitate controllable content generation in games [3, 60, 9]. World models are the primary tools for achieving this goal: they aim to capture an environment’s dynamics, predict its future states, and simulate its evolution [68, 29, 63, 50].
Most current world models focus on open-world settings [60, 35, 4, 8], generating unconstrained and often random environments in which each simulation produces a different world. This approach is intended to enhance generalization by exposing agents to diverse training scenarios [39, 3, 32, 31, 9]. However, the emphasis on randomness can lead to unstable dynamics, making such models ill-suited for applications that demand reliable prediction and precise planning in fixed tasks. In these cases, they often produce futures that are merely plausible, not faithful.
@@
  • Representation learning. Exteroceptive sensory data, such as images, are high-dimensional and encode complex, nonlinear mappings from underlying physical processes [5, 77, 70, 82, 44]. This makes accurate future-state prediction difficult even under full observability [29, 31]. For example, a simple spatial translation in the physical world can correspond to a highly nonlinear trajectory in pixel space [69, 24, 6]. Addressing this requires a representation space that faithfully encodes the underlying physical states while minimizing information loss and noise, thereby simplifying the subsequent dynamics modeling task [76, 22, 64, 45, 5, 89].
@@
  • Dynamics modeling. Even with an optimal representation, the dynamics model must capture a broad range of transition patterns — including 3D transformations, logical rules, causal dependencies, and temporal memory [80, 47, 61, 1, 33, 14, 38]. The challenge lies in building a unified mechanism that can accurately model all these regularities [80, 1, 51, 78].
@@
These challenges are inherently coupled: a poor representation forces the dynamics model to operate in a noisy, entangled latent space, increasing prediction complexity and reducing generalization; conversely, a well-structured representation is of limited utility if the dynamics model cannot capture the full spectrum of transition patterns [85, 20, 70, 7, 49, 90]. Progress toward generic neural world simulation therefore demands a co-design of representation learning and dynamics modeling, ensuring that the latent space is both physically meaningful and optimally aligned with the predictive capabilities of the dynamics model [29, 64, 32, 51, 15].
@@
In this work, we focus on the faithful cloning of deterministic environments — settings governed by fixed rules, such as a 3D maze with a static map. In such cases, the objective is not to produce a plausible world, but to reproduce the unique, true trajectory of the environment. We target these environments for three reasons: (1) Their consistent dynamics make them amenable to precise predictive modeling. (2) The absence of stochasticity eliminates uncertainty, enabling rigorous evaluation of a world model’s fidelity. (3) Many important real-world applications — including robot navigation in fixed spaces and game AI operating on static maps — inherently involve deterministic environments. More importantly, our primary goal is not merely to solve the cloning task itself, but to use deterministic cloning as a controlled setting for understanding the fundamental factors that limit long-horizon fidelity in world models.
@@
Yet, even in this controlled setting, we find that achieving accurate long-horizon cloning of simple deterministic environments remains an open challenge. Across all state-of-the-art baselines we evaluated, none were able to maintain fidelity over extended horizons: small prediction errors accumulate rapidly, causing trajectories to diverge from reality after only a few steps. In contrast, when the dynamics model is provided with the environment’s underlying physical states [^1] We refer to this version as the oracle model, since it has access to the true underlying states of the environment. — rather than high-dimensional exteroceptive signals such as images — it can produce remarkably accurate long-horizon predictions. This observation suggests that the effectiveness of a world model is fundamentally constrained by the structure of its latent representation space, and thus raises a natural question: How can we build a self-supervised representation that aligns with the underlying physical states, enabling stable and accurate long-horizon predictions?
@@
A natural next step is to investigate how to bridge the gap to oracle-level performance through unsupervised representation learning. We show that applying the temporal contrastive learning principle [23, 79, 59, 40, 19, 48] as a geometric regularization can effectively curate a latent space that better reflects the underlying physical state manifold, demonstrating that contrastive constraints can serve as a powerful inductive bias for stable world modeling; we call this approach Geometrically-Regularized World Models (GRWM). At its core is a lightweight geometric regularization module that can be seamlessly integrated into standard autoencoders, reshaping their latent space to better align with the underlying physical states and provide a stable foundation for effective dynamics modeling. Through extensive experiments, we demonstrate that improved representation geometry directly translates into higher long-horizon fidelity across multiple world models, validating the “representation matters” hypothesis. GRWM acts as a plug-and-play component that systematically unlocks long-horizon fidelity for state-of-the-art world models.
@@ 2 Related Work
Video world models aim to predict future visual observations of an environment conditioned on agent actions, often using latent generative modeling as a central approach [60, 35, 56]. Latent generative world models [64, 62, 29] first encode observations into a latent space via an autoencoder and then model transitions in this space to predict future frames. This paradigm is widely applicable [60, 56, 4, 9, 8] and has inspired recent works exploring various enhancements, including interactive controllability [53, 60, 12], architectural innovations for causality [84, 14] and persistence [88, 2, 28, 87, 66], and improved physical accuracy [13, 27, 52]. Despite these advances, the predominant focus has been on improving dynamics modeling, e.g., designing more expressive latent transition models [58, 38, 66, 2] and adding causal and relational inductive biases [55, 33, 14]. The autoencoder’s role is largely engineering: to provide a compressive latent embedding for the dynamics model [^2] Commonly, the autoencoder is trained with a small KL weight and additional perceptual or adversarial losses [62]., with most work focusing on improving reconstruction quality rather than shaping the latent geometry for stable long-horizon prediction.
Moreover, most video world models target open-world or stochastic environments, emphasizing diversity and plausibility over faithful reproduction of a unique trajectory [60, 35, 4, 9, 8]. In contrast, deterministic environments, where the goal is to precisely clone the true sequence of states, have received less attention. Our work highlights that in such settings, representation quality can be the primary bottleneck: even strong dynamics models fail if the latent space is poorly structured, motivating methods like GRWM that explicitly regularize latent geometry.
@@
Pure reconstruction objectives often yield degenerate representations that ignore temporal cues. Contrastive learning principles [30, 72, 17, 16, 34, 77, 83, 25, 75] suggest a remedy: enforce similarity for related samples and repulsion for unrelated ones. Several methods extend this to temporal data, such as Contrastive Learning Through Time (CLTT) [21, 65, 59, 81, 40, 41, 19, 86, 75], where contrastive objectives are used to capture temporal coherence in sequential observations.
@@ 3 Preliminary
The Next-State Prediction Problem. World models aim to perform next-state prediction: learning a function that predicts future observations given the current state and an action [10, 29, 67]. An environment evolves through latent states $s_{t}$, which yield observations $o_{t}$ after actions $a_{t}$. In the partially observable case, $o_{t}$ gives only partial information about $s_{t}$, requiring integration of past history to form a belief state. In the fully observable case, $s_{t}\approx o_{t}$, yet predicting in observation space remains difficult because simple physical transitions (e.g., translation) correspond to complex, non-linear changes in high-dimensional pixel space. This necessitates a learned representation space that can linearize these dynamics.
@@
We adopt a latent generative framework [64, 62, 29] consisting of two stages: (1) an autoencoder that learns a compressed latent representation [46, 36, 73], and (2) a generative model that captures the dynamics in that latent space [29, 31, 9]. Our work focuses on the representation stage, as its quality critically determines the downstream rollout performance.
@@ 4 Representation as the Bottleneck
Figure 1: Representation quality is the primary bottleneck for world model fidelity. Frame-wise MSE on a simple deterministic 3D navigation environment. (Left) An oracle model using ground-truth states (black dotted) achieves near-zero error, establishing a performance upper bound. In contrast, a standard VAE-based world model (blue dashed) accumulates error rapidly. Our GRWM (green solid) significantly closes this gap by learning a more structurally aligned latent space (Right Bottom), while the VAE’s representation remains disorganized (Right Top). For further details, see Section 6.
We begin with a diagnostic question: when a world model fails over long horizons, is the main limitation the dynamics model, or the representation on which that dynamics model operates? In this section, we use the term bottleneck to refer to the component that primarily limits rollout fidelity.
This question is especially important in visual world modeling. Even in a simple deterministic environment, the underlying state may be low-dimensional, while the model only observes high-dimensional pixel inputs. In that case, the model must first recover a useful state representation from images before it can predict future observations accurately. If this representation does not preserve the structure of the true state space, then small prediction errors can quickly accumulate during rollout.
To isolate the role of representation, we construct an “oracle” world model that bypasses the perception problem and operates directly on the ground-truth state variables, here given by the agent pose $(x,y,\theta)$. The oracle contains three components: (1) an encoder that maps a history of observations to the previous state: $\hat{s}_{t-1}=f_{\text{enc}}(o_{t-k:t-1}),$ (2) a dynamics model that predicts the current state from the previous state and action: $\hat{s}_{t}=f_{\text{dyn}}(\hat{s}_{t-1},a_{t}),$ and (3) a decoder that reconstructs the current observation from the predicted state: $\hat{o}_{t}=f_{\text{dec}}(\hat{s}_{t}).$
Figure 1 (left) shows the key result. The plotted MSE is pixel-space frame-wise error between predicted and ground-truth observations. Under this evaluation, the oracle maintains near-zero error over long horizons, while VAE-WM accumulates error rapidly. Since the two models are architecturally identical and differ only in the source of the latent state, this comparison leads to two conclusions: first, high-fidelity long-horizon cloning is feasible in this setting; second, the main limitation is not the dynamics model itself, but the representation on which it operates.
Why does a standard latent representation fail? A likely reason is geometric mismatch. A latent space trained mainly for reconstruction does not necessarily preserve the neighborhood structure or global layout of the true state space. As a result, points that are nearby in the true dynamics may not be nearby in latent space, and vice versa. This makes even simple dynamics harder to model reliably over long horizons.
Figure 1 (right) provides an intuitive illustration of this issue. The latent space learned by a standard VAE appears poorly organized with respect to the underlying state structure. This observation motivates the next section, where we introduce geometric regularization to encourage representations that are better aligned with the environment’s state manifold.
@@ 5 Geometrically-Regularized World Models
We next explore how to reduce the gap to oracle-level performance through unsupervised representation learning. Our approach applies the temporal contrastive learning principle [59, 40, 19, 48] as a geometric regularizer on the autoencoder of a world model. This regularizer shapes the latent space so that its geometry better matches the underlying state manifold, providing a strong inductive bias for stable long-horizon prediction.
@@ 5.1 Temporal Contextualize Architecture
Learning underlying states from a single frame is difficult due to perceptual aliasing [18, 54, 11, 43], where distinct states can yield nearly identical observations. For example, two different positions in a maze may appear visually indistinguishable. This also clarifies a key difference in the oracle experiment: the oracle uses coordinates, which encode the global state directly, whereas images provide only local information. A single image cannot specify the global position in environments with heavy aliasing. To recover the true state, the model must aggregate information over time, since a sequence of observations can form a unique signature of the underlying global state [59, 40, 19].
@@
Each frame is first encoded independently with a 2D CNN. The resulting frame-level features are then aggregated by a causal Transformer [74] with a sliding temporal window, ensuring that $z_{t}$ only attends to a limited range of past frames up to and including time $t$. This windowed design captures short-term temporal context while maintaining causality.